IMAGING OF ELECTRONICALLY INELASTIC COLLISIONS.

\underline{BRADLEY F. PARSONS}, AND DAVID W. CHANDLER, Combustion Research Facility, Sandia National Laboratory, 7011 East Ave., Livermore, CA 94550.

We have investigated the elastic and inelastic scattering in the Cl + D_{2} system as a probe of the coupling between Cl (^{2}P_{3/2}) + D_{2} potential energy surface and the Cl* (^{2}P_{1/2}) + D_{2} potential energy surface. Our experiments utilize a newly constructed cross molecular beam machine with velocity map ion imaging detection to probe all scattering into the Cl* (^{2}P_{1/2}) channel at a collision energy of 1700 cm^{−1}. Collisions between Cl (^{2}P_{1/2}) which result in a non-adiabatic transition to the upper spin-orbit potential energy surface must partition 881 cm^{−1} (=E_{S.O.,Cl}) less energy into relative translation. Our ion images show no evidence of such spin-orbit changing collisions in the Cl + D_{2} system and allow for an estimate of the total cross section for such collisions.